OpenStudy (cenaida):

The sum of five times a number and twice a second number is 18, but the difference of three times the first number and seven times the second number is -22. Find the two numbers using the elimination method. A. -2 and -4 B. -2 and 4 C. 2 and -4 D. 2 and 4

OpenStudy (anonymous):

5x+2y=18 3x-7y=-22

OpenStudy (anonymous):

let the two no be x and y 5x+2y=18 3x-7y=-22 can you solve these equations ?